Hello:
I am trying to answer the following math problem, but I do not know where to start:
The number of bacteria in a dish has tripled every month. If there were 2500 bacteria at first, how many bacteria were there 10 months later.
I have tried to answer this problem a couple ways:
using A=p^rt and by creating an equation. However, I am just not sure.
Please give me some guidance.
Thank you
LK

Answer by stanbon(48569)

The number of bacteria in a dish has tripled every month. If there were 2500 bacteria at first, how many bacteria were there 10 months later.
----------------
A(t) = Ao(3)^t
A(10) = 2500(3)^10
A(10) = 2500*59049
A(10) = 147,633,500
-----------
Cheers,
Stan H.